A force of 50 N was used to push a 10 kg object which was initially at rest.

a.) What is the resulting acceleration of the object as a result of that action of the force?

b.) What is the velocity now of the object 10 seconds after the force was first applied.
1
Expert's answer
2018-08-28T09:50:08-0400
(a) We can find the resulting acceleration of the object from the Newton’s Second Law of Motion:
F=ma,
a=F/m=(50 N)/(10 kg)=5 m/s^2 .
(b) We can find the velocity of the object 10 seconds after the force was first applied from the kinematic equation:
v=v_0+at,
here, v_0=0 is the initial velocity of the object, v is the final velocity of the object, a is the acceleration of the object and t is the time.
Then, we get:
v=v_0+at=5 m/s^2 ∙10 s=50 m/s.